# FAST Polarization Catalog of FRB 20240114A

### 21 Mar 2026 v1

@@ Abstract
Polarization measurements of fast radio bursts (FRBs) probe the magnetized plasma surrounding their central engines. FRB 20240114A is an exceptionally active repeating source, with 17,356 bursts detected between 2024 January 28 and 2025 May 30 by FAST, enabling time-resolved polarimetric studies. In this work, we present a polarimetric catalog of 6,131 bright bursts (with a signal-to-noise ratio S/N 20, 35.3% of the total sample), including arrival time (MJD), dispersion measure (DM), burst width (W), bandwidth, Faraday rotation measure (RM), linear and circular polarization degrees (DOL, DOC), and intrinsic polarization angle (PA0). We detect a clear temporal evolution of RM: after an initial stable phase, it decreases linearly by 200 over 200 days, forming a bimodal distribution, whereas DM remains stable at 528.9 . The linear polarization fraction is generally high, with the 3 lower bound around 76%, while circular polarization is low, with 1,157 of 17,356 bursts (6.67%) having DOC 10%. We perform a power-law fit between /I and , which yields an index of . It is found that the combined 2D distribution of L/I versus V/I remains stable, implying that the emission mechanism is largely invariant. Our PA0 measurements show a broad, non-uniform distribution, implying a complex emission geometry. These results suggest that FRB 20240114A resides in a dynamically evolving magneto-ionic environment. This catalog provides a foundation for studies of repeating FRB progenitors and their environments.
@@ I Introduction
Fast Radio Bursts (FRBs) represent one of the most energetic and enigmatic phenomena in the transient universe (Zhang, 2023). Since their discovery (Lorimer et al., 2007; Thornton et al., 2013), the identification of repeating sources (Spitler et al., 2016; CHIME/FRB Collaboration et al., 2019) has provided a crucial path toward understanding their progenitors. Among the various observational tools, radio polarimetry has proven to be exceptionally powerful. Measurements of the linear polarization and the Faraday rotation measure (RM) offer a direct probe into the magneto-ionic environment surrounding the FRB source, potentially revealing the presence of a supernova remnant, a pulsar wind nebula, or a binary companion’s wind (Michilli et al., 2018; Yang et al., 2020; Qu and Zhang, 2023). Furthermore, the temporal evolution of these properties can shed light on the dynamics of this local environment (Hilmarsson et al., 2021; Xu et al., 2022).
@@
The repeating FRB 20240114A, discovered shortly after its initial activity, has established itself as one of the most prolific known sources (Shin et al., 2025; Tian et al., 2024; Panda et al., 2025; Xie et al., 2025; Zhang et al., 2025a; Zhou et al., 2025b; Zhang et al., 2025b). Its high burst rate makes it an ideal target for amassing a large catalog of bursts. The Five-hundred-meter Aperture Spherical radio Telescope (FAST, Nan et al., 2011), with its unparalleled sensitivity, is uniquely capable of detecting the faint end of the burst energy distribution with a fluence completeness threshold of 0.026 Jy ms from such a source (Zhang et al., 2025a), thereby providing a complete and unbiased sample. This extreme sensitivity also makes FAST a powerful platform for high-precision polarimetry, enabling it to probe the extreme and dynamic magneto-ionic environments of repeating fast radio bursts, as demonstrated by a series of landmark studies (Luo et al., 2020; Xu et al., 2022; Jiang et al., 2022; Anna-Thomas et al., 2023; Niu et al., 2024; Jiang et al., 2024; Zhou et al., 2025a; Feng et al., 2025; Li et al., 2026).
This paper is the first in a series devoted to unveiling the polarization nature of FRB 20240114A through FAST observations. Herein, we focus exclusively on the presentation and public release of the polarization catalog. We describe our 16-month observations from 2024 to 2025, the detailed data reduction pipeline, and the calibration procedures that underpin this dataset. The resulting catalog comprises over 6000 bursts, each with measured arrival time (MJD), dispersion measure (DM), RM, and linear and circular polarization fractions.
@@ II.1 FAST Observations and Burst Search
An observation campaign of FRB 20240114A was conducted using the FAST telescope for over sixteen months, from 2024 January 28 to 2025 May 30. The initial observations employed the position reported by CHIME (Shin et al., 2025) (RA = 21h27m39s.888, Dec = +04∘21′0′′.36). Starting from 15 February 2024 (UT), the pointing was updated to the refined location (RA = 21h27m39s.84, Dec = +04∘19′46′′.3) localized by MeerKAT (Tian et al., 2024). The observing campaign comprised 97 sessions, totaling 57.99 hours of exposure time. Observations covered the 1.0–1.5 GHz band using the central beam of the 19-beam receiver in the tracking mode. Most data were recorded in the 8-bit PSRFITS format, providing full Stokes polarization information, 4096 frequency channels with a time resolution of 49.152 s.
We performed single-pulse searches on the FAST data using the PRESTO software package (Ransom, 2001), which employs a matched-filtering technique that applies boxcar functions of varying widths to the dedispersed time series to identify pulses across a range of timescales. The pipeline was described in Li et al. (2025). The search results underwent cross-verification using the developed deep learning-based search tool DRAFTS (Zhang et al., 2025c) and the AI-aided data filtering pipeline detailed in Wang et al. (2024). This process identified 17,356 bursts exceeding our detection threshold (fluence signal-to-noise ratio SNRf 12; see Zhang et al. (2025a)). From this sample, we selected 6,131 bursts with S/N 20 for polarization analysis.
@@ II.2 Parameter Measurement
The time of arrival (TOA) for each burst was determined by measuring the time at which the cumulative fluence reached 50% of its total value, following the burst identification criteria defined in Li et al. (2021) and Zhang et al. (2025b), which classify multiple pulses as components of a single burst when a complex time-frequency structure is bridged by emission above 5. The final TOA is taken as the topocentric arrival time (MJD) at the 1.5 GHz reference.
@@
For bursts exhibiting ambiguous morphology in the dynamic spectrum or relatively low S/N, the multi-component optimization may become unstable and lead to overfitting. In such cases, we adopt the daily averaged DM values to ensure consistency across the sample. The derived DMs have a median value of 529.3 with a standard deviation of 1.2 , consistent with the measurement reported by CHIME (Shin et al., 2025).
The burst frequency boundaries were determined using the CDF-based method proposed by Xie et al. (2025). For each burst, the spectrum was first constructed by integrating the on-pulse flux density over time for each frequency channel. The cumulative distribution function (CDF) of the spectrum was then calculated and smoothed using the asymmetrically reweighted penalized least squares (arPLS) algorithm to suppress noise fluctuations. The lower and upper frequency limits were defined as the frequencies where the first derivative of the normalized CDF rises above and subsequently falls below thresholds determined by the signal-to-noise ratio (S/N), specifically 1/SNR for the rising edge and 1–1/SNR for the falling edge. This procedure identifies the dominant emission band based on significant spectral power gradients rather than assuming a Gaussian spectral shape, and is therefore not equivalent to Gaussian fitting.
@@
We employed the RM synthesis method to search for the rotation measure of each burst within a range of -20000 to 20000 , successfully measuring RM for 6,131 bursts. To ensure data reliability, we examined each burst for potential saturation effects in the FAST receiver. By inspecting the four polarization components (XX, YY, X∗Y, Y∗X) and identifying data points at the critical thresholds, we flagged and subsequently excluded bursts exhibiting genuine saturation signatures. This screening removed 253 bursts from the initial sample with S/N20, ensuring that our polarization analysis is not affected by non-linear instrumental effects (see Appendix A for details).
@@
We additionally conducted a systematic search for Faraday conversion across all bursts. Faraday conversion is a propagation effect in a magnetized plasma with a perpendicular magnetic field, converting linear polarization into circular polarization. Combined with Faraday rotation, if the Faraday conversion is observed, the detection would constrain plasma properties and magnetic field geometry along the line of sight (Wang et al., 2025). Our analysis found no significant evidence of conversion in any of the bursts in our sample, consistent with the results of Uttarkar et al. (2026), indicating that the observed circular polarization is likely intrinsic to the emission mechanism rather than being generated by propagation effects.
@@ III The Catalog Description
| L/I | % | Debiased linear polarization fraction. |
| V/I | % | Circular polarization fraction. |
@@
| Burst ID | MJD | DM () | RM () | DOL (%) | DOC (%) | Weff (ms) | Bandwidth (MHz) | S/N |
| B00001 | 60337.23827975 | 529.50.1 | 328.5 | 96.32.1 | -7.11.5 | 3.510.23 | 188.310.6 | 29.7 |
| B00007 | 60337.25012753 | 527.70.1 | 317.5 | 97.32.8 | -12.32.0 | 3.020.21 | 207.76.2 | 38.3 |
| B00009 | 60338.21737789 | 528.50.1 | 326.5 | 97.13.1 | -26.12.3 | 1.620.06 | 319.28.1 | 46.9 |
| B00011 | 60338.22518088 | 527.20.6 | 335.5 | 60.43.8 | -82.44.2 | 2.520.14 | 74.56.7 | 20.4 |
| B00012 | 60338.22802442 | 528.30.1 | 330.5 | 98.82.1 | 1.01.5 | 2.00.07 | 158.721.2 | 42.8 |
| B00018 | 60338.23511480 | 531.10.6 | 330.5 | 102.32.8 | -11.61.9 | 3.960.2 | 155.54.0 | 22.0 |
| B00019 | 60338.23556485 | 529.10.1 | 327.5 | 92.12.6 | -0.71.9 | 2.470.1 | 174.121.5 | 33.0 |
| B00028 | 60341.19895450 | 530.40.5 | 343.5 | 54.62.5 | 11.22.2 | 3.650.2 | 121.36.6 | 20.5 |
| B00031 | 60349.15748811 | 529.00.3 | 348.5 | 98.41.8 | 18.41.3 | 1.940.09 | 90.214.3 | 28.8 |
| B00032 | 60349.15783813 | 528.50.1 | 350.5 | 96.11.2 | -6.50.9 | 4.910.16 | 237.69.4 | 34.6 |
| B14288 | 60824.89193507 | 530.90.1 | 246.1 | 99.32.2 | -5.11.5 | 3.090.02 | 276.45.9 | 56.0 |
| B14292 | 60824.89230168 | 530.00.1 | 231.9 | 86.39.4 | 1.87.1 | 2.810.03 | 325.112.6 | 57.4 |
| B14293 | 60824.89238065 | 529.80.0 | 250.7 | 99.22.2 | -10.81.6 | 1.980.01 | 458.99.4 | 152.0 |
| B14297 | 60824.89327416 | 529.60.2 | 248.1 | 103.83.0 | 1.42.1 | 3.990.02 | 442.514.7 | 54.7 |
| B14298 | 60824.89356171 | 530.20.1 | 242.7 | 91.51.9 | -22.91.5 | 3.040.03 | 204.110.3 | 53.6 |
| B14302 | 60824.89386021 | 529.80.1 | 245.3 | 97.70.7 | 9.90.5 | 3.650.01 | 255.58.8 | 177.9 |
| B14304 | 60824.89416654 | 529.40.2 | 243.5 | 100.21.5 | -2.91.0 | 2.550.02 | 195.66.7 | 106.0 |
| B14313 | 60824.89625329 | 530.60.1 | 240.1 | 90.98.7 | 16.16.5 | 2.190.07 | 287.84.2 | 26.4 |
| B14314 | 60824.89625000 | 530.50.1 | 239.5 | 94.44.7 | 0.93.4 | 2.120.05 | 213.08.6 | 40.8 |

### 18 Apr 2026 v2

@@ Abstract
Polarization measurements of fast radio bursts (FRBs) probe the magnetized plasma surrounding their central engines. FRB 20240114A is an exceptionally active repeating source, with 17,356 bursts detected between 2024 January 28 and 2025 May 30 by FAST, enabling studies of the temporal evolution of its polarization properties. In this work, we present a polarimetric catalog of 6,131 bright bursts (with a signal-to-noise ratio S/N 20, 35.3% of the total sample), including arrival time (MJD), dispersion measure (DM), burst width (W), bandwidth, Faraday rotation measure (RM), linear and circular polarization degrees (, ), and intrinsic polarization angle (PA0). We confirm a clear temporal evolution of RM: after an initial stable phase, it decreases linearly by 200 over 200 days, forming a bimodal distribution, whereas DM remains stable at 529.3 1.2 . The linear polarization fraction is generally high, with the 3 lower bound around 76%, while circular polarization is low, with 1,157 of 17,356 bursts (6.67%) having 10%. We perform a power-law fit between and , which yields an index of . It is found that the combined 2D distribution of versus remains stable, implying that the emission mechanism is largely invariant. Our PA0 measurements show a broad, non-uniform distribution, implying a complex emission geometry. These results suggest that FRB 20240114A resides in a dynamically evolving magneto-ionic environment. This catalog provides a foundation for studies of repeating FRB progenitors and their environments.
@@ I Introduction
Fast radio bursts (FRBs) represent one of the most energetic and enigmatic phenomena in the transient universe (Zhang, 2023). Since their discovery (Lorimer et al., 2007; Thornton et al., 2013), the identification of repeating sources (Spitler et al., 2016; CHIME/FRB Collaboration et al., 2019) has provided a crucial path toward understanding their progenitors. Among the various observational tools, radio polarimetry has proven to be exceptionally powerful. Measurements of the linear polarization and the Faraday rotation measure (RM) offer a direct probe into the magneto-ionic environment surrounding the FRB source, potentially revealing the presence of a supernova remnant, a pulsar wind nebula, or a binary companion’s wind (Michilli et al., 2018; Yang et al., 2020; Wang et al., 2022; Qu and Zhang, 2023). Furthermore, the temporal evolution of these properties can shed light on the dynamics of this local environment (Hilmarsson et al., 2021; Xu et al., 2022).
@@
Recent years have also seen the emergence of systematic FRB polarization catalogs from multiple instruments. For non-repeating FRBs, polarimetric catalogs have been presented for samples detected with DSA-110 (Sherman et al., 2024), CHIME/FRB (Pandhi et al., 2024), and ASKAP (Scott et al., 2025). For repeating sources, recent CHIME/FRB studies have established the largest source-level polarization catalogs to date (Mckinven et al., 2023c; Ng et al., 2025), enabling comparative studies across repeater populations. In this context, the present work is complementary: rather than a source-level census, we provide a large burst-level, time-resolved polarization catalog for a single exceptionally active repeating FRB, making it possible to investigate burst-to-burst variations and long-term temporal evolution within one source.
The repeating FRB 20240114A, discovered shortly after its initial activity, has established itself as one of the most prolific known sources (Shin et al., 2025; Tian et al., 2024; Panda et al., 2025; Xie et al., 2025; Zhang et al., 2025a; Zhou et al., 2025b; Zhang et al., 2025b).Its sustained high burst rate over an extended period makes it an ideal target for amassing a large catalog of bursts. The Five-hundred-meter Aperture Spherical radio Telescope (FAST, Nan et al., 2011), with its unparalleled sensitivity, is uniquely capable of detecting the faint end of the burst energy distribution with a fluence completeness threshold of 0.026 Jy ms from such a source (Zhang et al., 2025a), thereby providing a complete and unbiased sample. This extreme sensitivity also makes FAST a powerful platform for high-precision polarimetry. Previous FAST observations have revealed diverse polarization-angle swings, frequency-dependent polarization behavior, extreme circular polarization, and temporal evolution of the local magneto-ionic environment in repeating FRBs (Luo et al., 2020; Xu et al., 2022; Jiang et al., 2022; Anna-Thomas et al., 2023a; Niu et al., 2024; Jiang et al., 2024; Zhou et al., 2025a; Feng et al., 2025; Li et al., 2026), demonstrating that polarimetric measurements can serve as a sensitive probe of their extreme and dynamic environments. Building on these results, the present work provides a systematic polarimetric catalog and time-resolved statistical analysis of FRB 20240114A based on a much larger burst sample.
This paper is the first in a series devoted to unveiling the polarization nature of FRB 20240114A through FAST observations. Herein, we focus exclusively on the presentation and public release of the polarization catalog. We describe our 16-month observations from 2024 to 2025, the detailed data reduction pipeline, and the calibration procedures that underpin this dataset. The resulting catalog comprises over 6000 bursts, each with measured time of arrival (TOA), dispersion measure (DM), RM, and linear and circular polarization fractions.
@@ II.1 FAST Observations and Burst Search
An observation campaign of FRB 20240114A was conducted using the FAST telescope for over sixteen months, from 2024 January 28 to 2025 May 30. The initial observations employed the position reported by the Canadian Hydrogen Intensity Mapping Experiment (CHIME) (Shin et al., 2025) (RA = 21h27m39s.888, Dec = +04∘21′0′′.36). Starting from 15 February 2024 (UT), the pointing was updated to the refined location (RA = 21h27m39s.84, Dec = +04∘19′46′′.3) localized by MeerKAT (Tian et al., 2024). The observing campaign comprised 97 sessions, totaling 57.99 hours of exposure time. Observations covered the 1.0–1.5 GHz band using the central beam of the 19-beam receiver in the tracking mode. Most data were recorded in the 8-bit PSRFITS format, providing full Stokes polarization information, 4096 frequency channels with a frequency resolution of 0.1221 MHz and a time resolution of 49.152 s.
We performed single-pulse searches on the FAST data using the PRESTO software package (Ransom, 2001), which employs a matched-filtering technique that applies boxcar functions of varying widths to the dedispersed time series to identify pulses across a range of timescales. The pipeline was described in Li et al. (2025). To evaluate the completeness of the burst search, we also cross-checked the PRESTO results using two additional search approaches: DRAFTS, a deep-learning-based radio fast transient search pipeline (Zhang et al., 2025c), and the AI-aided pipeline described in Wang et al. (2024). In the present work, PRESTO serves as the primary burst-search pipeline, while these two methods are used as auxiliary cross-verification tools rather than as independent parallel catalog-building pipelines. By comparing their candidate lists with the PRESTO detections on the same data set, we verified that there is no obvious population of missed bursts in the final sample. This process identified 17,356 bursts exceeding our detection threshold, defined by a fluence signal-to-noise ratio of SNRf 12, where SNRf denotes the signal-to-noise ratio of the burst fluence, i.e., the integrated burst flux over its duration, rather than the peak-flux signal-to-noise ratio (see Zhang et al., 2025a). From this sample, we selected 6,384 bursts with S/N 20 for polarization analysis.
@@ II.2 Parameter Measurement
The time of arrival (TOA) for each burst was determined by measuring the time at which the cumulative fluence reached 50% of its total value, following the burst identification criteria defined in Li et al. (2021) and Zhang et al. (2025b), which classify multiple pulses as components of a single burst when a complex time-frequency structure is bridged by emission above 5. The final TOA is taken as the topocentric arrival time (MJD) at the 1.5 GHz reference. For the polarization analysis, the effective temporal limits of each burst are defined by its measured TOA and full width at half maximum (FWHM), adopting a time window of TOAFWHM. The FWHM is obtained by fitting a Gaussian to the autocorrelation function of the same profile.
@@
For bursts exhibiting ambiguous morphology in the dynamic spectrum or relatively low S/N, the multi-component optimization may become unstable and lead to overfitting. In such cases, we adopt the daily averaged DM values to ensure consistency across the sample. The derived DMs have a median value of 529.3 with a standard deviation of 1.2 , consistent with the measurement reported by CHIME/FRB (Shin et al., 2025).
The burst frequency boundaries were determined using the CDF-based method proposed by Xie et al. (2025). For each burst, the spectrum was first constructed by integrating the on-pulse flux density over time for each frequency channel. The cumulative distribution function (CDF) of the spectrum was then calculated and smoothed using the asymmetrically reweighted penalized least squares (arPLS) algorithm to suppress noise fluctuations. The lower and upper frequency limits were defined as the frequencies where the first derivative of the normalized CDF rises above and subsequently falls below thresholds determined by the signal-to-noise ratio (S/N), specifically for the rising edge and for the falling edge. This procedure identifies the dominant emission band based on significant spectral power gradients rather than assuming a Gaussian spectral shape, and is therefore not equivalent to Gaussian fitting.
@@
We applied the RM synthesis method to all 6,384 bursts with S/N20, searching for rotation measures over the range from to (Brentjens and de Bruyn, 2005). The adopted range was chosen as a conservative symmetric interval, comfortably wider than the RMs actually measured for FRB 20240114A. For the present FAST L-band channelization, intrachannel depolarization would become important only at much larger , and therefore does not limit the current RM measurements. To ensure data reliability, we examined each burst for potential saturation effects in the FAST receiver. By inspecting the four polarization components (XX, YY, X∗Y, Y∗X) and identifying data points at the critical thresholds, we flagged and subsequently excluded bursts exhibiting genuine saturation signatures. This screening removed 253 bursts from the initial sample with S/N20, leaving a final sample of 6,131 nonsaturated bursts with reliable RM measurements for the polarization analysis (see Appendix A for details).
@@
We additionally conducted a systematic search for Faraday conversion across all bursts. Faraday conversion is a propagation effect in a magnetized plasma with a perpendicular magnetic field, converting linear polarization into circular polarization. Combined with Faraday rotation, if the Faraday conversion is observed, the detection would constrain plasma properties and magnetic field geometry along the line of sight (Wang et al., 2025). Our analysis found no significant evidence of conversion in any of the bursts in our sample, consistent with the results of Uttarkar et al. (2026b), indicating that the observed circular polarization is likely intrinsic to the emission mechanism rather than being generated by propagation effects.
@@ III The Catalog Description
For the convenience of catalog users, we also provide estimates of the Galactic foreground contributions toward FRB 20240114A. Using the YMW16 and NE2025 Galactic electron-density models (Yao et al., 2017; Ocker and Cordes, 2026), we obtain Galactic interstellar dispersion measures of DMMW,ISM = 38.824.78 pc cm-3 and 51.057.45 pc cm-3, respectively. The quoted uncertainties correspond to fractional model uncertainties of 12.32% for YMW16 and 14.59% for NE2025, estimated from pulsars with measured DMs at similar Galactic latitude by comparing the observed DMs with the corresponding model predictions. Using the YT2020 Galactic halo model (Yamasaki and Totani, 2020), we further estimate a halo contribution of DMMW,halo = 45.05 pc cm-3, where the uncertainty is obtained by applying the 0.2 dex (rms) fluctuation reported for the Milky Way halo DM in that model. We also estimate the Galactic Faraday rotation foreground from the Galactic Faraday rotation sky map (Hutschenreuter et al., 2022), obtaining RMMW = -14.55 10.48 rad m-2. These values are not burst-by-burst catalog entries, but source-level line-of-sight foreground references that may be useful for interpreting the measured DMs and RMs in the catalog.
@@
| % | Debiased linear polarization fraction. | |
| % | Circular polarization fraction. | |
@@
| Burst ID | MJD | DM | RM | Weff | Bandwidth | PA0 | S/N | ||
| () | () | (%) | (%) | (ms) | (MHz) | (deg) | |||
| B00001 | 60337.23827975 | 529.50.1 | 328.5 | 96.32.1 | -7.11.5 | 3.510.23 | 188.310.6 | -7.8 | 29.7 |
| B00007 | 60337.25012753 | 527.70.1 | 317.5 | 97.32.8 | -12.32.0 | 3.020.21 | 207.76.2 | 46.3 | 38.3 |
| B00009 | 60338.21737789 | 528.50.1 | 326.5 | 97.13.1 | -26.12.3 | 1.620.06 | 319.28.1 | 19.0 | 46.9 |
| B00011 | 60338.22518088 | 527.20.6 | 335.5 | 60.43.8 | -82.44.2 | 2.520.14 | 74.56.7 | 7.3 | 20.4 |
| B00012 | 60338.22802442 | 528.30.1 | 330.5 | 98.82.1 | 1.01.5 | 2.00.07 | 158.721.2 | 73.1 | 42.8 |
| B00018 | 60338.23511480 | 531.10.6 | 330.5 | 102.32.8 | -11.61.9 | 3.960.2 | 155.54.0 | 11.8 | 22.0 |
| B00019 | 60338.23556485 | 529.10.1 | 327.5 | 92.12.6 | -0.71.9 | 2.470.1 | 174.121.5 | 8.7 | 33.0 |
| B00028 | 60341.19895450 | 530.40.5 | 343.5 | 54.62.5 | 11.22.2 | 3.650.2 | 121.36.6 | -53.4 | 20.5 |
| B00031 | 60349.15748811 | 529.00.3 | 348.5 | 98.41.8 | 18.41.3 | 1.940.09 | 90.214.3 | -48.8 | 28.8 |
| B00032 | 60349.15783813 | 528.50.1 | 350.5 | 96.11.2 | -6.50.9 | 4.910.16 | 237.69.4 | 23.6 | 34.6 |
| B17290 | 60824.89193507 | 530.90.1 | 246.1 | 99.32.2 | -5.11.5 | 3.090.02 | 276.45.9 | 7.4 | 56.0 |
| B17310 | 60824.89230168 | 530.00.1 | 231.9 | 86.39.4 | 1.87.1 | 2.810.03 | 325.112.6 | 52.6 | 57.4 |
| B17313 | 60824.89238065 | 529.80.1 | 250.7 | 99.22.2 | -10.81.6 | 1.980.01 | 458.99.4 | -50.2 | 152.0 |
| B17315 | 60824.89327416 | 529.60.2 | 248.1 | 103.83.0 | 1.42.1 | 3.990.02 | 442.514.7 | -11.2 | 54.7 |
| B17317 | 60824.89356171 | 530.20.1 | 242.7 | 91.51.9 | -22.91.5 | 3.040.03 | 204.110.3 | 35.1 | 53.6 |
| B17318 | 60824.89386021 | 529.80.1 | 245.3 | 97.70.7 | 9.90.5 | 3.650.01 | 255.58.8 | 59.2 | 177.9 |
| B17328 | 60824.89416654 | 529.40.2 | 243.5 | 100.21.5 | -2.91.0 | 2.550.02 | 195.66.7 | -64.3 | 106.0 |
| B17332 | 60824.89625329 | 530.60.1 | 240.1 | 90.98.7 | 16.16.5 | 2.190.07 | 287.84.2 | 11.0 | 26.4 |
| B17344 | 60824.89625000 | 530.50.1 | 239.5 | 94.44.7 | 0.93.4 | 2.120.05 | 213.08.6 | 33.1 | 40.8 |
Note. All uncertainties in this table are rounded to one decimal place, following Zhang et al. (2025a) for consistency with previously published results. For a small subset of bursts whose DM uncertainties would be rounded to 0.0 under this convention, we uniformly adopt a conservative floor value of 0.1 pc cm-3to avoid reporting unphysical zero-valued uncertainties. The quoted RM uncertainties are empirical estimates based on the RM-synthesis peak width and burst S/N. An additional systematic uncertainty comparable to one RM-search grid step ( 0.2 rad m-2) may also be considered.